Hunter & Sons sells a single model of meat smoker for use in the home. The smokers have the following price and cost characteristics.

Sales price $ 84 per smoker
Variable costs 32 per smoker
Fixed costs 395,200 per month

Hunter & Sons is subject to an income tax rate of 40 percent.

Required:

a. How many smokers must Hunter & Sons sell every month to break even?

b. How many smokers must Hunter & Sons sell to earn a monthly operating profit of $106,080 after taxes?

a. Break-even sales in units b. Number of units to be sold
Answer #1

a)Breakeven point is a situation of No profit no loss that is Total revenue is equals to total cost

Breakeven point in units = Fixed cost /(price-variable cost)

                        = 395200 /(84 -32)

                        = 395200 / 52

                        = 7600 units

b)Desired profit before tax =After tax profit /(1-tax rate)

                          = 106080 /(1-.40)

                           = 106080 /.60

                             = $ 176800

Number of units to sold to ear desired income =[Fixed cost+Before tax income ]/[sales price- variable cost]

                         = [ 395200+176800] /(84 -32)

                         = 572000/ 52

                          = 11000units